OpenStudy (anonymous):

find the sum ∑_(n=1)^100▒2n

OpenStudy (anonymous):

hmmm

OpenStudy (australopithecus):

is this summation censored?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[\sum_{n=1}^{100}2n\]?

OpenStudy (anonymous):

yes

OpenStudy (kinggeorge):

\[\sum_1^{100} 2n=2 \cdot \sum_1^{100} n=2\cdot {n(n+1) \over 2}\]

OpenStudy (kinggeorge):

Where \(n=100\)

OpenStudy (anonymous):

\[\sum_{k=1}^nk=\frac{n(n+1)}{2}\] so ... what king george said

OpenStudy (anonymous):

still not getting it

OpenStudy (australopithecus):

you can add them all up like 2(1) + 2(2) + 2(3)... or you can just use a formula to solve for it which is much quicker

OpenStudy (anonymous):

we are counting from 1 to 100 what is the nth term.

OpenStudy (australopithecus):

100 terms sorry

OpenStudy (anonymous):

2xn(n+1)/2 so i use this fromula to solve it.

OpenStudy (anonymous):

in general \[\sum_{i=1}^{n} i={n(n+1)\over2}\]

OpenStudy (australopithecus):

(2)100(100+1)/2 = 10100 proof https://www.wolframalpha.com/input/?i=summation+of+1+to+100+of+2n

OpenStudy (australopithecus):

if you want to understand the formula I recommend just looking up sum of arithmetic formula on wiki

OpenStudy (australopithecus):

or you can just accept it

OpenStudy (anonymous):

the rules of summations are such that we can take constants out of the summation sign\[\sum_{i=1}^{n} Ci=C\sum_{i=1}^{n} i=C{n(n+1)\over2}\]where C is any constant that is how we got our formula

OpenStudy (anonymous):

The proof is quite easy if you know mathematical induction if not, it's better to just accept it, or look up the formula for any arithmetic series the wiki way, as suggested by @Australopithecus

OpenStudy (anonymous):

okay so c=2 n=100

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay got it thanks

OpenStudy (anonymous):

welcome